Income Statement Balance Sheet
  Sales $ 96,600   Current assets $ 36,500   Debt $ 47,400
  Costs 68,850   Fixed assets 95,900   Equity 85,000
  Taxable income $ 27,750      Total $ 132,400      Total $ 132,400  
  Tax 6,938
  Net Income $ 20,812

Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 40 percent dividend payout ratio. No external financing is possible.

What is the internal growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Internal growth rate= ROA × b / 1-ROA × b

b= retention ratio= 1- dividend payout ratio

b=1-0.4=0.6

Return on Assets = Net income / total Assets

ROA= 20812/ 132400= 15.72%

Internal growth rate= (0.1572×0.6) /(1-(0.1572×0.6))

Internal growth rate= 0.09432/ (1-0.09432)= 10.41%